# Preference-Guided Reinforcement Learning for Efficient Exploration

### v1

@@ C-A LOPE Policy Architecture for Discrete Control Tasks
LOPE Policy Architecture for Continuous Control Tasks
@@
LOPE Value Function Architecture
@@ C-B Hyper-parameter settings
The hyper-parameters settings are presented in Table I.
@@ Appendix D Additional Implementation Details
For clarity in the presentation, we omit certain implementation details about experiment environment settings in Section VI and describe them here. The locomotion control and ant maze tasks were implemented based on the rllab framework (https://github.com/rll/rllab), a framework used to develop and evaluate reinforcement learning algorithms. We ran the Hopper and HalfCheetah agents in a straight channel. Their observations consisted of the agent’s joint angle information and task-specific attributes. These agents can receive sparse rewards by reaching the specified locations. We set the length of the channel to 10 units for SparseHopper and 90 units for SparseHalfCheetah. These robots were asked to move as far forward as possible. Meanwhile, the ant robot was required to walk through the E-shaped maze, as shown in Fig. 2d. This maze is a benchmark environment for hierarchical reinforcement learning [56].

### v2

@@ C-A Neural Network Architectures
LOPE policy architecture for continuous control and navigation tasks:
@@
LOPE value function architecture:
@@ C-B Hyper-parameter settings
The hyperparameter settings for different tasks are presented in Table I.
@@ Appendix D Additional Implementation Details
For clarity in the presentation, we omit certain implementation details about experiment environment settings in Section VI and describe them here. The locomotion control and maze navigation tasks were implemented using MuJoCo [todorov2012mujoco] and D4RL [fu2020d4rl], which are widely used for developing and evaluating reinforcement learning algorithms. We ran the Hopper and HalfCheetah agents in a straight channel. Their observations consisted of the agent’s joint angle information. These agents can receive sparse rewards by reaching the specified locations. We set the length of the channel to 10 units for SparseHopper and 90 units for SparseHalfCheetah. These robots were asked to move as far forward as possible. In the AntMaze-Umaze task, the ant robot was required to walk through the U-shaped maze, as shown in Fig. 3b. Its observations consisted of the agent’s joint angle information and task-specific attributes. This maze is a benchmarking test environment for maze navigation problems. In the PointMaze tasks, a 2-DoF ball is force-actuated in the Cartesian directions x and y to reach a target goal in a closed maze.
Appendix E Additional experimental results
E-A MMD distance changing trend
We investigate the changing trend of the MMD distance between the agent’s policy $\pi$ and the preferred trajectories in $\mathcal{P}$. Fig. 13a shows the changing trend of the MMD distance between $\pi$ and $\mathcal{P}$ during the training process of LOPE in the grid-world task. In the early stage of the training process, a considerable MMD distance was maintained between $\pi$ and $\pi_{b}$. Then, the MMD distance dropped rapidly. Fig. 13b shows the changing trend of the MMD distance between $\pi$ and $\pi_{b}$ during the training process of LOPE in the AntMaze task. The MMD distance decreases gradually during training. Intuitively, the policy $\pi_{b}$ implied by the preference dataset $\mathcal{P}$ is superior to the current policy $\pi$. Updating the policy parameters towards $\pi_{b}$ guarantees a higher performance improvement bound.
Figure 13: Changing trends of MMD distance in the fixed grid-world maze and ant maze.